The measure of an angle is 6 more than twice the measure of its complement. Find the measures of both angles. Thank you.
Answer by askmemath(368)

Complementary angles sum up to 90degrees.
Let the smaller angle be X
Then the larger angle becomes 2X+6
Sum = X + 2X+6 = 3X+6
3X +6 = 90
3X = 84
X = 84/3 = 28
Larger angle = 2(28)+6 = 62